Please help me solve this equation:
2/ (x+5) + 1/ (x-5) = 16/ (x^2 -25)
I know the answer equals 7, but I am unable to retrieve that answer.
Here is my work:
LCD= (x+5)(x-5)
I multiplied the LCD with the equation and cancelled out the bottom fractions.
2x - 10 + x +5 = 0
3x - 5 = 0
3x = 5
x = 5/3
Thanks for your help!

Answer by mananth(16946)

2/ (x+5) + 1/ (x-5) = 16/ (x^2 -25)
You have equated to 0 . instead of 16
2(x-5)+(x+5)=16
2x-10+x+5=16
3x=16+10-5
3x=21
/3
x=7